# Tri-resonant leptogenesis from modular symmetry neutrino models

### v1

@@ 2.2 Some basics about RGE effects
$T_{IJ}$ is singular at the scale where $M_{J}=M_{I}$ holds unless ${\rm Re}[(YY^{\dagger})_{IJ}]=0$. To achieve this condition, we perform an additional rotation in the above basis:
$M^{\prime}_{\rm R}=O^{\dagger}M_{\rm R}O^{*}=M_{\rm R}\;,\qquad Y^{\prime}=O^{\dagger}Y\;,$ (22)
where $O$ is a real orthogonal matrix which can be parameterized by three angles $\theta_{1,2,3}$. This rotation does not affect the right-handed neutrino mass matrix due to their exact mass degeneracy. $YY^{\dagger}$ becomes
$Y^{\prime}Y^{\prime\dagger}=O^{\dagger}YY^{\dagger}O\;.$ (23)
Apparently, the three condition ${\rm Re}[(YY^{\dagger})_{IJ}]=0$ can be always satisfied by choosing suitable vuales of $\theta_{1,2,3}$. The RGE evolution of the right-handed neutrino mass matrix in the new basis is
$4\pi^{2}\frac{{\rm d}M_{I}}{{\rm d}t}=\left(Y^{\prime}Y^{\prime\dagger}\right)_{II}M_{I}\;.$ (24)
For illustration, taking account of the RGE effects from the flavor-symmetry scale $\Lambda_{\rm FS}$ to the right-handed neutrino mass scale $M_{0}$, one can obtain the following relation for the right-handed neutrino masses
$M_{I}\left(M_{0}\right)\simeq M_{I}\left(\Lambda_{\rm FS}\right)\left[1-\frac{1}{4\pi^{2}}\left(Y^{\prime}Y^{\prime\dagger}\right)_{II}\ln\left(\frac{\Lambda_{\rm FS}}{M_{0}}\right)\right]\;,$ (25)
yielding the following mass splittings for the right-handed neutrinos
$M_{I}\left(M_{0}\right)-M_{J}\left(M_{0}\right)\simeq\frac{M_{I}\left(\Lambda_{\rm FS}\right)}{4\pi^{2}}\left[\left(Y^{\prime}Y^{\prime\dagger}\right)_{JJ}-\left(Y^{\prime}Y^{\prime\dagger}\right)_{II}\right]\ln\left(\frac{\Lambda_{\rm FS}}{M_{0}}\right)\;.$ (26)
Such mass splittings among the right-handed neutrinos are comparable to their decay widths
$\Gamma_{I}=\frac{\left(YY^{\dagger}\right)_{II}}{8\pi}M_{I}\left(M_{0}\right)\;,$ (27)
fulfilling the condition for the resonant leptogenesis to proceed. Note that in obtaining Eq. (25), as a rough estimate, we have only considered the diagonal-term [i.e., $(YY^{\dagger})_{II}$-related] contributions to the right-handed neutrino masses. But the off-diagonal-term [i.e., $(YY^{\dagger})_{IJ}$-related] contributions may be equally important in the case of the right-handed neutrino masses being degenerate, so in the following calculations we will also consider them.
@@ 3 Study for modular ${\rm A}_{4}$ symmetry model
while the right-handed neutrino mass matrix takes the same form as in Eq. (8). On the other hand, in the case that three right-handed charged leptons are successively assigned to the singlet representations ${\bf 1}$, ${\bf 1^{\prime\prime}}$ and ${\bf 1^{\prime}}$ of $A_{4}$, the charged-lepton mass matrix takes a form as
@@
with $\omega=e^{{\rm i}2\pi/3}$. One can see that this model totally contains 8 real parameters: ${\rm Re}(\tau)$, ${\rm Im}(\tau)$, $\beta/\alpha$, $\gamma/\alpha$, $|g_{2}/g_{1}|$, ${\rm arg}(g_{2}/g_{1})$, $\alpha v_{d}$ and $g^{2}_{1}v^{2}_{u}/M_{0}$. The three parameters $\alpha v_{d}$, $\beta/\alpha$ and $\gamma/\alpha$ can be fixed by fitting with the values of three charged lepton masses. The remaining parameters serve to describe the three neutrino masses, three neutrino mixing angles and three CP phases present in the neutrino mixing matrix. The detailed numerical calculation shows that this model can be consistent with the neutrino oscillation experimental results in both the NO and IO cases. And the resulting predictions for the best-fit values and the allowed ranges of the input parameters and observables are collected in Table 2.
@@
${\rm Im}(\tau)$ 1.824 $[0.91,1.16]\cup[1.31,1.86]$ 0.987 $[0.98,1.049]\cup[1.052,1.109]$
$\beta/\alpha$ 205.720 $[192.39,215]\cup[3054.254093.49]$ 79.472 $[59.68,86.37]\cup[892.67,1446.02]$
$\gamma/\alpha$ 3612.07 $[192.4,215]\cup[3066.47,4092.98]$ 1232.57 $[60.97,86.34]\cup[870.16,1443.84]$
$|g_{2}/g_{1}|$ 2.410 $[2.398,2.71]\cup[2.95,3.86]$ 2.093 $[1.038,2.453]$
$\arg{(g_{2}/g_{1})}$ 6.267 $[0,0.49]\cup[6.23,2\pi]$ 4.715 $[1.33,1.83]\cup[4.29,5]$
@@
Table 2: For the modular ${\rm A}_{4}$ symmetry model given in section 3, the predictions for the best-fit values and the allowed ranges of the input parameters and observables in the NO and IO cases. These results are taken from Tables 6 and 10 in Ref. [20].
We first consider the scenario that the mass splittings among the three right-handed neutrinos are realized by modifying $M_{\rm R}$ into the form as shown in Eq. (10). In order to facilitate the leptogenesis calculation, we go to the basis where both the right-handed neutrino mass matrix and charged lepton mass matrix are diagonal with the help of $U_{\rm R}$ in Eq. (9) (which serves to diagonalize the the right-handed neutrino mass matrix) and another unitary matrix $U_{\rm L}$ which serves to diagonalize the charged lepton mass matrix. At the same time, the neutrino Yukawa coupling matrix becomes $Y^{\prime}=U^{\dagger}_{\rm R}YU_{\rm L}$ which is relevant for the leptogenesis calculations. With help of the Eq. (13) and the fitted values of the input parameters as given in Table 2, we numerically calculate the CP asymmetries in the decays of $N_{I}$ as functions of $2\mu/\Gamma_{I}$. The results are shown in Figure 1. Since $\mu$ is a free parameter, the resonance conditions $2\mu/\Gamma_{I}\simeq 1$ which enhance the CP asymmetries to their maximally allowed values can always be fulfilled. This will be helpful for us to reproduce the observed value of $\eta_{\rm B}$. Figure 2 has further shown the ratios of various collision terms in the Boltzmann equations to the Hubble constant as functions of $z=M_{0}/T$ with $M_{0}=10^{3}$ GeV and $\mu=10^{-8}$ GeV as benchmark values (see below for the reason why we have taken $\mu=10^{-8}$ GeV as a benchmark value). We see that the scattering collision term $\Gamma^{S}_{\rm Gauge}$ is comparable to $\Gamma^{S}_{\rm Yukawa}$, and they provide the dominant sources of the scattering collision terms in the temperature range $T>M_{0}$. On the other hand, in the temperature range $T<M_{0}$, the dominant term is $\Gamma^{D}_{\rm Yukawa}$ that describes the decay process. Then, the numerical scan finds the values of $\mu$ that accommodate a successful leptogenesis as functions of $M_{0}$ as shown in Figure 3. Considering that people are more interested in low-scale leptogenesis (which has the potential to be directly accessed by running or upcoming experiments), our scan over $M_{0}$ is performed in the range from $10^{3}$ GeV to $10^{6}$ GeV. The results show that for each value of $M_{0}$ there exist two values of $\mu$ that accommodate a successful leptogenesis, and the ratio of $\mu$ to $M_{0}$ is about $10^{-10}$ or $10^{-14}$ ($10^{-9}$ or $10^{-21}$) in the NO (IO) cases. For the purpose of illustration, by taking $M_{0}=10^{3}$ GeV as a benchmark value, in Figure 4 we have shown the evolutions of $\eta_{N}$ and $\eta_{B}$ as functions of $z$. These results are obtained by taking the initial conditions to be $\eta^{\rm ini}_{N}=\eta^{\rm ini}_{B}=0$. The results show that, for the benchmark value $M_{0}=10^{3}$ GeV, the observed value of $\eta_{\rm B}$ can be successfully reproduced for $\mu\simeq 10^{-8.4}$ GeV ($10^{-5.9}$ GeV) in the NO (IO) case.
Figure 1: For the modular ${\rm A}_{4}$ symmetry model given in section 3, in the scenario that the mass splittings among the three right-handed neutrinos are realized by modifying $M_{\rm R}$ into the form as shown in Eq. (10), CP asymmetries in the decays of $N_{I}$ as functions of the ratio $2\mu/\Gamma_{I}$ in the NO (left panel) and IO (right panel) cases. Here $\varepsilon_{\rm tot}=\varepsilon_{1}+\varepsilon_{2}+\varepsilon_{3}$ has been defined.
Figure 2: For the modular ${\rm A}_{4}$ symmetry model given in section 3, in the scenario that the mass splittings among the three right-handed neutrinos are realized by modifying $M_{\rm R}$ into the form as shown in Eq. (10), the various collision terms in the Boltzmann equations as functions of $z=M_{0}/T$ in the NO (left panel) and IO (right panel) cases. These results are obtained by taking $M_{0}=10^{3}$ GeV and $\mu=10^{-8}$ GeV as benchmark values.
Figure 3: For the modular ${\rm A}_{4}$ symmetry model given in section 3, in the scenario that the mass splittings among the three right-handed neutrinos are realized by modifying $M_{\rm R}$ into the form as shown in Eq. (10), the values of $\mu$ that accommodate a successful leptogenesis as functions of $M_{0}$ in the NO (left panel) and IO (right panel) cases.
Figure 4: For the modular ${\rm A}_{4}$ symmetry model given in section 3, in the scenario that the mass splittings among the three right-handed neutrinos are realized by modifying $M_{\rm R}$ into the form as shown in Eq. (10), the evolutions of $\eta_{N}$ and $\eta_{B}$ as functions of $z$ for the benchmark value $M_{0}=10^{3}$ GeV and some benchmark values of $\mu$ in the NO (left panel) and IO (right panel) cases. Here the horizontal grey dashed line shows the observed value of $\eta_{\rm B}$ (same as below) and the vertical dotted line represents the decoupling temperature $T_{*}\simeq 131.7$ GeV of sphalerons [21].
Then, we consider the scenario that the mass splittings among the three right-handed neutrinos are generated from the RGE effects. In this regard, Figure 5 has shown the RGE-induced mass splittings of the three right-handed neutrinos. In obtaining these results, we have fixed the flavor-symmetry scale $\Lambda_{\rm FS}$ at the grand unification (GUT) scale $10^{16}$ GeV. The results show that the ratios $\Delta M_{IJ}/M_{0}$ (for $\Delta M_{IJ}\equiv M_{I}-M_{J}$) are linearly proportional to $M_{0}$. This is because $\Delta M_{IJ}/M_{0}$ are proportional to $YY^{\dagger}$ [see Eq. (26)], which in turn is proportional to $M_{0}$ because of the seesaw relationship [see Eq. (6)]. Figure 6 further shows the CP asymmetries in the decays of $N_{I}$ as functions of $M_{0}$. In the NO case, for $M_{0}<10^{11}$ GeV, the CP asymmetries are small and nearly independent of $M_{0}$. But for $M_{0}>10^{11}$ GeV, $\varepsilon_{3}$ increases rapidly with $M_{0}$, while $\varepsilon_{1}$ and $\varepsilon_{2}$ also do so after a sign reversal from being positive to negative. In the IO case, the CP asymmetries keep positive and change very slowly as $M_{0}$ increases. Finally, Figure 7 has shown the generated $\eta_{B}$ as functions of $M_{0}$. As one can see, the observed value of $\eta_{B}$ can be successfully reproduced for $M_{0}\simeq 6\times 10^{12}$ GeV ($4\times 10^{12}$ GeV or $6\times 10^{12}$ GeV) in the NO (IO) case. These results can be understood with the help of the results in Figure 6.
Figure 5: For the modular ${\rm A}_{4}$ symmetry model given in section 3, in the scenario that the mass splittings among the three right-handed neutrinos are generated from the RGE effects, $\Delta M_{IJ}/M_{0}$ (for $\Delta M_{IJ}\equiv M_{I}-M_{J}$) as functions of $M_{0}$ in the NO (left panel) and IO (right panel) cases.
@@
Figure 6: For the modular ${\rm A}_{4}$ symmetry model given in section 3, in the scenario that the mass splittings among the three right-handed neutrinos are generated from the RGE effects, the CP asymmetries in the decays of $N_{I}$ as functions of $M_{0}$ in the NO (left panel) and IO (right panel) cases.
Figure 7: For the modular ${\rm A}_{4}$ symmetry model given in section 3, in the scenario that the mass splittings among the three right-handed neutrinos are generated from the RGE effects, the generated $\eta_{\rm B}$ as functions of $M_{0}$ in the NO (in the red color) and IO (in the blue color) cases.

### v3

@@ 2.2 Some formulas about RGE effects
It is immediate to see that $T_{IJ}$ is singular at the scale where one has $M_{J}=M_{I}$, unless ${\rm Re}[(YY^{\dagger})_{IJ}]=0$ holds. Fortunately, due to the exact mass degeneracy of three right-handed neutrinos, one has a freedom of rotation $N\to ON$ with $O^{T}O=I$ for them, which does not affect the unity right-handed neutrino mass matrix, but rotates the Yukawa coupling matrix to the following form
$Y\to Y^{\prime}=OY\;.$ (30)
Such a freedom can be utilized to assure that Eq. (28) is non-singular [20], namely,
${\rm Re}\left[Y^{\prime}Y^{\prime\dagger}\right]={\rm Re}\left[O^{\dagger}YY^{\dagger}O\right]=0\;.$ (31)
3 Study for a model realizing the TM1 mixing
In this section, we perform the study for a neutrino flavor-symmetry model realizing the popular TM1 mixing [21]. To be specific, we will consider the concrete model presented in section 3 of Ref. [22], which is constructed based on the ${\rm S}_{4}$ flavor symmetry in the MSSM framework. ${\rm S}_{4}$ is the permutation group of 4 objects with 24 elements, and it has two singlet representations $\bf 1$} and $\bf 1^{\prime}$}, one two-dimensional representation $\bf 2$} and two three-dimensional representations $\bf 3$} and $\bf 3^{\prime}$}. In this model both the three left-handed lepton doublets and three right-handed neutrinos constitute a $\bf 3^{\prime}$} representation of ${\rm S}_{4}$, while three right-handed charged leptons are trivially $\bf 1$} representations of ${\rm S}_{4}$. In order to be able to yield certain predictions for the CP phases at the same time, the flavor symmetry is extended in such a way that a generalized CP symmetry [26] is also included. As is commonly done in the paradigm of flavor symmetries, the full symmetries including both the flavor and CP symmetries are broken (by the flavon VEVs) into different subgroups in the neutrino and charged-lepton sectors, and the mismatch between them gives rise to certain predictions for the lepton mixing angles and CP phases.
In the model under consideration, thanks to the residual flavor and CP symmetries, the charged-lepton mass matrix is constrained to be diagonal, and the Dirac neutrino mass matrix is constrained into the following form
$M_{\rm D}=\frac{y_{1}v_{u}v_{\xi}}{\Lambda_{\rm FS}}\left[\begin{pmatrix}1&0&0\\ 0&0&1\\ 0&1&0\end{pmatrix}+x\begin{pmatrix}0&1&1\\ 1&1&0\\ 1&0&1\end{pmatrix}+{\rm i}y\begin{pmatrix}0&1&-1\\ -1&0&-2\\ 1&2&0\end{pmatrix}\right$]\;, (32)
while the right-handed neutrino mass matrix takes the same form as in Eq. (15). Here $v_{u}$ and $v_{\xi}$ are the VEVs of the up-type Higgs and flavon fields, $y_{1}$, $x$ and $y$ are real parameters. Given the above forms of $M_{\rm D}$ and $M_{\rm R}$, the seesaw formula yields a light neutrino mass matrix as
$M_{\nu}=a\left(\begin{matrix}2&-1&-1\cr-1&2&-1\cr-1&-1&2\end{matrix}\right)+b\left(\begin{matrix}1&0&0\cr 0&0&1\cr 0&1&0\end{matrix}\right)+c\left(\begin{matrix}0&1&1\cr 1&1&0\cr 1&0&1\end{matrix}\right)+{\rm i}d\left(\begin{matrix}0&1&-1\cr 1&2&0\cr-1&0&-2\end{matrix}\right)\;,$ (33)
with
$a=m_{0}y^{2}\;,\hskip 28.45274ptb=-m_{0}(1+2x^{2}+4y^{2})\;,$
$c=-m_{0}(2x+x^{2}+y^{2})\;,\hskip 28.45274ptd=-3m_{0}xy\;,$ (34)
and $m_{0}=y^{2}_{1}v^{2}_{u}v^{2}_{\xi}/(M_{0}\Lambda^{2}_{\rm FS})$. It is easy to verify that such an $M_{\nu}$ not only keeps invariant under the TM1 transformation but also under the $\mu$-$\tau$ reflection transformation [27]:
$M_{\nu}=R^{T}_{\rm TM1}M_{\nu}R_{\rm TM1}\hskip 14.22636pt{\rm with}\hskip 14.22636ptR_{\rm TM1}=-\frac{1}{3}\left(\begin{array}[]{ccc}1&2&2\cr 2&-2&1\cr 2&1&-2\cr\end{array}\right)\;,$ (38)
$M_{\nu}=R^{T}_{\mu\tau}M^{*}_{\nu}R_{\mu\tau}\hskip 14.22636pt{\rm with}\hskip 14.22636ptR_{\mu\tau}=\left(\begin{array}[]{ccc}1&0&0\cr 0&0&1\cr 0&1&0\cr\end{array}\right)\;.$ (42)
Correspondingly, the resulting neutrino mixing not only has the features of the TM1 mixing but also the features of the $\mu$-$\tau$ reflection symmetry:
$\theta_{23}=\frac{\pi}{4}\;,\hskip 28.45274pt\delta=\pm\frac{\pi}{2}\;,\hskip 28.45274pt\alpha_{21}=0\ {\rm or}\ \frac{\pi}{2}\;,\hskip 28.45274pt\alpha_{31}=0\ {\rm or}\ \frac{\pi}{2}\;.$ (43)
Given that the considered model is subject to further constraints on top of the TM1 mixing, if it can successfully reproduce the observed value of $Y_{\rm B}$ with the help of the RGE effects, then the models that are only subject to the TM1 mixing necessarily can also do so.
It is found that this model can only be phenomenologically viable in the NO case and the best-fit values of $x$ and $y$ are x=0.050x=-0.050 and $y=\pm 0.233$ (corresponding to $\delta=\pm\pi/2$) [22]. And the parameter combination $y_{1}v_{u}v_{\xi}/\Lambda_{\rm FS}$ in $M_{\rm D}$ can be expressed as $\sqrt{m_{ee}M_{0}/(1+2x^{2}+2y^{2})}$ with $m_{ee}\simeq 57.9$ meV (i.e., the effective neutrino mass for neutrinoless double beta decays). The most stringent bounds on neutrinoless double beta decay lifetimes come from the KamLAND-Zen [23] and GERDA [24] experiments, which constrain the effective Majorana mass as $m_{ee}<$ (28-122)meV and $m_{ee}<$ (79-180) meV, respectively. The model also predicts the masses of the three neutrinos: $m_{1}=57.284$ meV, $m_{2}=57.930$ meV and $m_{3}=75.488$ meV. For comparison, the latest Planck results constrain the sum of neutrino masses to $\sum_{i}m_{i}<$ (0.12-0.60)eV [25]. Furthermore, the Yukawa coupling matrix can be obtained as $Y=M_{\rm D}/v_{u}$ with $v_{u}=v\tan\beta/\sqrt{1+\tan^{2}\beta}$. Here $\tan\beta$ is the ratio of the VEV of the up-type Higgs field $H_{u}$ to that of the down-type Higgs field $H_{d}$. Note that in the MSSM a moderately large $\tan\beta$ value (e.g., 30) is favored, for which the relation $v_{u}\simeq v$ holds.
Now, we are ready to study the RGE induced tri-resonant leptogenesis for the above model: at the flavor-symmetry scale $\Lambda_{\rm FS}$, three right-handed neutrinos are completely degenerate, prohibiting leptogenesis to proceed; if there is a considerable gap between $\Lambda_{\rm FS}$ and the leptogenesis scale (roughly $M_{0}$), then the RGE effects from the former scale to the latter scale will lift the degeneracy among three right-handed neutrino masses, inducing a tri-resonant leptogenesis to work. Given that most parameters of the model have been fixed as described above, there are only three free parameters relevant to our study: $M_{0}$, $\Lambda_{\rm FS}$ and $\tan\beta$. Considering that the RGE effects only depend on $\Lambda_{\rm FS}$ in a logarithmic manner (and we have numerically checked that the dependence of the final results on the concrete value of $\Lambda_{\rm FS}$ is weak), in the following numerical calculations we will take $\Lambda_{\rm FS}=10^{16}$ GeV (i.e., around the grand unification scale) as a benchmark value. Now that $M_{0}$ and $\tan\beta$ are the remaining free parameters (note that the value of $\tan\beta$ will directly affect the sizes of the charged-lepton Yukawa coupling and consequently the flavor effects), in Figure 1 we have shown the RGE-induced mass splittings of the three right-handed neutrinos as functions of $M_{0}$ for the benchmark values of $\tan\beta=10$ and 50. The results show that the ratios $\Delta M_{IJ}/M_{0}$ (for $\Delta M_{IJ}\equiv M_{I}-M_{J}$) are nearly linearly proportional to $M_{0}$ and insensitive to $\tan\beta$. This is because $\Delta M_{IJ}/M_{0}$ are proportional to $YY^{\dagger}$ (see Eq. (28)) which in turn is proportional to $M_{0}$. And $YY^{\dagger}$ is almost independent of $\tan\beta$ when it takes values from 10 to 50 (as mentioned in the end of last paragraph).
Figure 1: For the TM1 model in section 3, the mass splittings among the three right-handed neutrinos generated from the RGE effects, as functions of $M_{0}$ for the benchmark values of $\tan\beta=10$ and 50.
Figure 2: For the TM1 model in section 3, the resulting CP asymmetries for the decays of $N_{I}$ as functions of $M_{0}$ and $\tan\beta$.
Figure 2 has shown the resulting CP asymmetries $\varepsilon_{I}={\rm Tr}\;\varepsilon^{\left(I\right)}_{\alpha\beta}$ for the decays of $N_{I}$ as functions of $M_{0}$ and $\tan\beta$. The results show that $\varepsilon_{I}$ and $\varepsilon_{T}=\sum_{I}\varepsilon_{I}$ decrease with the increasing of $M_{0}$, and increase with $\tan\beta$. And the sign of $\varepsilon_{3}$ changes from positive to negative at $M_{0}\simeq 10^{12}$ GeV. Please keep in mind that here and in the following, for negative results, we have used the dashed lines to show their absolute values. Finally, Figure 3 has shown the resulting $\eta_{\rm B}$ as functions of $M_{0}$ for the benchmark values of $\tan\beta=10$, 20, 30 and 50. One can see that for $M_{0}>10^{12}$ GeV the four curves nearly converge into a single one. This means that in such a mass range the resulting $\eta_{B}$ is almost independent of $\tan\beta$. And the observed value of $\eta_{\rm B}$ can be successfully reproduced for $M_{0}\sim 7\times 10^{12}$ GeV.
Figure 3: For the TM1 model in section 3, the resulting $\eta_{\rm B}$ as functions of $M_{0}$ for some benchmark values of $\tan\beta$. The horizontal gray dashed line shows the observed value of $\eta_{\rm B}$.
@@ 4 Study for modular ${\rm A}_{4}$ symmetry model
while the right-handed neutrino mass matrix takes the same form as in Eq. (15). On the other hand, in the case that three right-handed charged leptons are successively assigned to the singlet representations $\bf 1$}, $\bf 1^{\prime\prime}$} and $\bf 1^{\prime}$} of ${\rm A}_{4}$, the charged-lepton mass matrix takes a form as
@@
with $\omega=e^{{\rm i}2\pi/3}$. One can see that this model totally contains 8 real parameters: ${\rm Re}(\tau)$, ${\rm Im}(\tau)$, $\beta/\alpha$, $\gamma/\alpha$, $|g_{2}/g_{1}|$, ${\rm arg}(g_{2}/g_{1})$, $\alpha v_{d}$ and $g^{2}_{1}v^{2}_{u}/M_{0}$. The three parameters $\alpha v_{d}$, $\beta/\alpha$ and $\gamma/\alpha$ can be fixed by fitting with the values of three charged lepton masses. The remaining parameters serve to describe the three neutrino masses, three neutrino mixing angles and three CP phases present in the neutrino mixing matrix. The detailed numerical calculation shows that this model can be consistent with the neutrino oscillation experimental results in both the NO and IO cases. And the resulting predictions for the best-fit values and the allowed ranges of the input parameters and observables are collected in Table 1 here.
@@
${\rm Im}(\tau)$ 1.824 $0.91,1.16]\cup[1.31,1.86]$ 0.987 $0.98,1.049]\cup[1.052,1.109]$
$\beta/\alpha$ 205.720 $192.39,215]\cup[3054.25,4093.49]$ 79.472 $59.68,86.37]\cup[892.67,1446.02]$
$\gamma/\alpha$ 3612.07 $192.4,215]\cup[3066.47,4092.98]$ 1232.57 $60.97,86.34]\cup[870.16,1443.84]$
$|g_{2}/g_{1}|$ 2.410 $2.398,2.71]\cup[2.95,3.86]$ 2.093 $[1.038,2.453]$
$\arg{(g_{2}/g_{1})}$ 6.267 $0,0.49]\cup[6.23,2\pi]$ 4.715 $1.33,1.83]\cup[4.29,5]$
@@
Table 1: For the modular ${\rm A}_{4}$ symmetry model in section 4, the predictions for the best-fit values and the allowed ranges of the input parameters and observables in the NO and IO cases. These results are taken from Tables 6 and 10 in Ref. [28].
Now, we study the RGE induced tri-resonant leptogenesis for the above model. For the purpose of illustration, Figure 4 has shown the CP asymmetries for the decays of $N_{I}$ as functions of $M_{0}$ for the benchmark value of $\tan\beta=30$. In the NO case, for $M_{0}<10^{12}$ GeV, the CP asymmetries are nearly independent of $M_{0}$. But for $M_{0}>10^{12}$ GeV, the absolute value of $\varepsilon_{3}$ increases with $M_{0}$, while $\varepsilon_{1}$, $\varepsilon_{2}$ and $\varepsilon_{T}$ also do so after a sign reversal from being positive to negative. In the IO case, the CP asymmetries keep their signs unchanged and change very slowly as $M_{0}$ increases. Figure 5 has further shown the CP asymmetries for the decays of $N_{I}$ as functions of $\tan\beta$ for the benchmark value of $M_{0}=10^{12}$ GeV. In both the NO and IO cases, the CP asymmetries increase with $\tan\beta$. Figure 6 has shown the resulting $\eta_{\rm B}$ as functions of $M_{0}$ for some benchmark values of $\tan\beta$. As one can see, in the NO case, the observed value of $\eta_{\rm B}$ can be successfully reproduced when $M_{0}$ is approximately between $10^{12}$ GeV and $10^{13}$ GeV. Figure 7 has further shown the values of $M_{0}$ versus $\tan\beta$ that allow for a successful reproduction of the observed value of $\eta_{\rm B}$. But in the IO case, the resulting $\eta_{\rm B}$ is always larger than the observed value, since the corresponding CP asymmetries are too large (see the right panel of Figure 4).
@@
Figure 4: For the modular ${\rm A}_{4}$ symmetry model in section 4, the resulting CP asymmetries for the decays of $N_{I}$ as functions of $M_{0}$ for the benchmark value of $\tan\beta=30$ in the NO (left panel) and IO (right panel) cases.
Figure 5: For the modular ${\rm A}_{4}$ symmetry model in section 4, the resulting CP asymmetries for the decays of $N_{I}$ as functions of $\tan\beta$ for the benchmark value of $M_{0}=10^{12}$ GeV in the NO (left panel) and IO (right panel) cases.
Figure 6: For the modular ${\rm A}_{4}$ symmetry model in section 4, the resulting $\eta_{\rm B}$ as functions of $M_{0}$ for the benchmark values of $\tan\beta=$10, 20, 30 and 50 in the NO (left panel) and IO (right panel) cases. The horizontal gray dashed line shows the observed value of $\eta_{\rm B}$.
Figure 7: For the modular ${\rm A}_{4}$ symmetry model in section 4, in the NO case, the values of $M_{0}$ versus $\tan\beta$ that allow for a successful reproduction of the observed value of $\eta_{\rm B}$.